# DNGaussian: Optimizing Sparse-View 3D Gaussian Radiance Fields with Global-Local Depth Normalization

### v1

@@ Abstract
Radiance fields have demonstrated impressive performance in synthesizing novel views from sparse input views, yet prevailing methods suffer from high training costs and slow inference speed. This paper introduces DNGaussian, a depth-regularized framework based on 3D Gaussian radiance fields, offering real-time and high-quality few-shot novel view synthesis at low costs. Our motivation stems from the highly efficient representation and surprising quality of the recent 3D Gaussian Splatting, despite it will encounter a geometry degradation when input views decrease. In the Gaussian radiance fields, we find this degradation in scene geometry primarily lined to the positioning of Gaussian primitives and can be mitigated by depth constraint. Consequently, we propose a Hard and Soft Depth Regularization to restore accurate scene geometry under coarse monocular depth supervision while maintaining a fine-grained color appearance. To further refine detailed geometry reshaping, we introduce Global-Local Depth Normalization, enhancing the focus on small local depth changes. Extensive experiments on LLFF, DTU, and Blender datasets demonstrate that DNGaussian outperforms state-of-the-art methods, achieving comparable or better results with significantly reduced memory cost, a $25\times$ reduction in training time, and over $3000\times$ faster rendering speed. Code is available at: https://github.com/Fictionarry/DNGaussian .
@@
Corresponding author: Xiao Bai (baixiao@buaa.edu.cn).
@@ 1 Introduction
Integrating the two proposed techniques, DNGaussian synthesizes views with competitive quality and superior details compared to state-of-the-art methods in multiple sparse-view settings on LLFF, Blender, and DTU datasets. This advantage is further enriched by substantially lower memory costs, $25\times$ reduction of training time, and over $3000\times$ faster rendering speed. The experiments also demonstrate our method’s universal ability to fit complex scenes, wide-ranging views, and multiple materials.
@@ 2 Related Work
Radiance Fields for Novel View Synthesis. Novel view synthesis aims to generate unseen views of the same object or scene from a set of given images [59, 1]. NeRF [25] is one of the most popular methods in recent years, which uses a large MLP to represent 3D scenes and renders via volume rendering. However, its speed is slow both in training and inference. The following improvements mainly pursue either higher quality [2, 3] or efficiency [6, 26, 37, 12, 53, 21, 15], but hard to achieve both. The most recent unstructured radiance fields [7, 51, 18] utilize a set of primitives to represent scenes. Among them, 3D Gaussian Splatting [18] represents radiance fields by a set of anisotropic 3D Gaussians and renders with a differentiable splatting. This approach achieves great success in fast and high-quality reconstruction for complex real scenes, and shows an advantage in high-frequency details. While this method excels with dense input views and has achieved success in various 3D tasks [23, 47, 39], its reconstruction with sparse view inputs remains an open problem. Also, issues such as how to apply additional constraints for improvement are still unsolved and worthy of discussion.
@@
Depth Supervision in Sparse-view Neural Fields. As a classic cue in many 3D vision tasks [46, 44, 41, 57, 43], depth information has been widely used to supervise sparse-view neural fields. The first group [11, 31] is to extract accurate but sparse depth values from reliable point clouds, and the second [55, 14, 36, 40, 42] distills depth knowledge from current powerful monocular depth estimators [30, 29]. Considering point clouds are sparse and not available in many sparse-view cases, monocular depth shows its advantage in density and robustness for our tasks. To tackle the scale ambiguity and error of monocular depths, previous works have introduced various scale-invariant losses [55, 10, 36] including depth ranking loss [50, 42], however, all of which are not optimal for us. These losses align the depth to a certain fixed global scale, which may ignore minor local depth changes. This overlook can lead to a noisy distribution of Gaussian primitives, particularly in regions with intricate textures. Besides, we notice an HDN loss [56] that can preserve depth details for monocular depth estimation. Nevertheless, this approach is also unsuitable as its reliance on multi-scale patches would bring long-distance errors and compromise geometric accuracy.
@@ 3.1 Preliminary for 3D Gaussian Splatting
Representation. 3D Gaussian splatting [18] represents 3D information with a set of 3D Gaussians. It computes pixel-wise color $\mathcal{C}$ with a set of 3D Gaussian primitives $\theta$, view pose $P$, and the camera parameter involving the center $\mathbf{o}$.
@@
$\mathcal{G}_{i}(\mathbf{x})=e^{-\frac{1}{2}(\mathbf{x-\mu_{i}})^{T}\Sigma_{i}^{-1}(\mathbf{x-\mu_{i}})},$ (1)
@@
Rendering. 3D Gaussian Splatting utilizes a point-based rendering to compute the color $\mathcal{C}$ of pixel $\mathbf{x}_{p}$ by blending $N$ ordered Gaussians overlapping the pixel:
@@
$\widetilde{\alpha}_{i}=\alpha_{i}\mathcal{G}^{proj}_{i}(\mathbf{x}_{p}).$ (3)
Then, similar to NeRF, we can represent the pixel-wise depth $\mathcal{D}$ with the distance to the camera center $\mathbf{o}$:
[…]
@@ 3.2 Depth Regularization for Gaussians
Hard Depth Regularization. To achieve the spatial reshaping of the Gaussian fields, we first propose a Hard Depth Regularization that encourages the movement of the nearest Gaussians, which are expected to compose surfaces but often cause noises and artifacts. Considering the predicted depth is rendered with the mixture of multiple Gaussians and reweighted by the cumulative product $\widetilde{\alpha}$, we manually apply a large opacity value $\tau$ to all Gaussians. Then, we render a ”hard depth” that mainly consists of the nearest Gaussians on the ray shot from camera center $\mathbf{o}$ and across the pixel $\mathbf{x}_{p}$:
[…]
@@
Soft Depth Regularization. Only regularizing on ”hard depth” is insufficient due to the absence of opacity optimization. We also expect to ensure the accuracy of the real rendered ”soft depth”, otherwise, the surface may become semitransparent and cause hollowness. From this perspective, we additionally freeze the Gaussian center $\mu$ (denoted by $\check{\mu}$) to avoid negative influence caused by the center moving, and propose Soft Depth Regularization for the tuning of the opacity $\alpha$:
[…]
@@ 3.3 Global-Local Depth Normalization
Local Depth Normalization. To solve this problem, we make the loss function refocus on small errors by introducing a patch-wise local normalization. Specifically, we cut a whole depth map into small patches and normalize the patch $\mathcal{P}$ of predicted depth and monocular depth with the mean value of 0 and standard deviation of near to $1$:
$\mathcal{D}^{LN}(\mathrm{x})=\frac{\mathcal{D}(\mathrm{x})-\text{mean}(\mathcal{D}(\mathcal{P}))}{\text{std}(\mathcal{D}(\mathcal{P}))+\epsilon},\quad\mathrm{s.t.}\ \ \mathrm{x}\in\mathcal{P},$ (8)
@@
$\mathcal{D}^{GN}(\mathrm{x})=\frac{\mathcal{D}(\mathrm{x})-\text{mean}(\mathcal{D}(\mathcal{P}))}{\text{std}(\mathcal{D}_{I})},$ (9)
$\mathrm{s.t.}\ \ \mathrm{x}\in\mathcal{P},$ $\mathcal{P}\subseteq\mathcal{I}.$
@@ 3.4 Training Details
[…]
@@
where $T$ stands for hardhard or softsoft. In practice, we reserve an error tolerance for the L2 loss to relax the constraint. The full loss function is formulated by:
@@
MVSNeRF ft [5]
Trained on DTU
Fine-tuned per Scene
17.88 0.327 0.584 0.157 18.54 0.197 0.769 0.113
@@
Neural Color Renderer. 3D Gaussian Splatting represents the color via spherical harmonic, however, it is easy to overfit with only sparse input views. To relieve this problem, we take a grid encoder and an MLP as the Neural Color Renderer to predict color for each primitive (Figure 3). During inference, we store the median result and only calculate the last MLP layers to merge view direction for acceleration.
@@ 4.1 Setups
Datasets. we conduct our experiment on three datasets: the NeRF Blender Synthetic dataset (Blender) [25], the DTU dataset [17], and the LLFF dataset [24]. We follow the setting used in previous works [27, 52, 42] with the same split of DTU and LLFF to train the model on 3 views and test on another set of images. To erase the noises in the background and focus on the target object, we apply the same object masks as previous works [27] for DTU at evaluation. For Blender, we follow DietNeRF [16] and FreeNeRF [52] to train with the same 8 views and test on 25 unseen images. Aligned with the baselines, downsampling rates of $8$, 44, and $2$ are applied to LLFF, DTU, and Blender. Following previous sparse-view settings, the camera poses are assumed to be known via calibration or other ways.
@@ 4.2 Comparison
Efficiency. We further conduct an efficiency study on the LLFF 3-view setting with RTX 3090 Ti GPUs to explore the performance of current SOTA baselines [42, 52] with limited GPU memories of 24GB/12GB, and training time of 1.0h/0.5h, as shown in Table 3. The top row of each group represents the default setting of the corresponding baseline, where the training time is measured by us with the same number of iterations on a single GPU. While both FreeNeRF and SparseNeRF perform worse under strict resource limitations, our method shows huge advantages in efficiency, which achieves remarkable accelerations of $25\times$ on training time and over $3000\times$ on FPS, while synthesizing competitive quality novel views. Given the necessity for per-scene optimization and rapid visualization, our high efficiency holds significant value for practical applications.
@@ 5 Conclusion
Acknowledgements. In this work, we are supported by the National Natural Science Foundation of China 62276016, 62372029. Lin Gu was supported by JST Moonshot R&D Grant Number JPMJMS2011 Japan.
@@ A.2 Ablation Study on Neural Color Renderer
In this work, we replace the spherical harmonic (SH) of 3D Gaussian Splatting with a neural color renderer to represent the direction-variant color. To better illustrate the function of this module, we compare it to the original SH function with different degrees in the LLFF dataset with 3 training views. The results are in Table 7 and Figure 11. The SH function is easy to overfit in the sparse-view situation and results in some strange colors during view changing. This may be caused by the independence of each primitive which leads to a lack of regional consistency. After introducing the neural color renderer, the problem has been relieved. By storing the median result and only calculating the latest two MLP layers, we can maintain a fast rendering speed competitive to SH as well.
@@ B.1 Implementations
Metrics. Following previous methods [27, 52], we utilize the ”structural_similarity” API from scikit-image  [^3] https://scikit-image.org/docs/stable/api/skimage.metrics.html to compute the SSIM score, and use the implementation with a pre-trained AlexNet model to calculate the LPIPS score.
@@ B.2 Datasets
LLFF. The LLFF dataset [24] contains 8 forward-facing scenes in total. Following [27, 52, 42], we take every 8-th image as the novel views for testing. The input views are evenly sampled across the remaining views. Images are downsampled $8\times$ to the resolution of $378\times 504$. In practice, we ignore the distortion of the original images.
DTU. The DTU dataset [17] consists of 124 object-centric scenes captured by a set of fixed cameras. We follow [27, 52, 42] to evaluate models directly on the 15 scenes with the scan IDs of 8, 21, 30, 31, 34, 38, 40, 41, 45, 55, 63, 82, 103, 110, and 114. In each scan, the images with the following IDs of 25, 22, and 28 are used as the input views in our 3-view setting. The test set consists of images with IDs of 1, 2, 9, 10, 11, 12, 14, 15, 23, 24, 26, 27, 29, 30, 31, 32, 33, 34, 35, 41, 42, 43, 45, 46 and 47 for evaluation. The images are downsampled $4\times$.
Blender. We follow the data split used in [16, 52] for the Blender dataset [11]. The 8 input views are selected from the training images, with IDs 26, 86, 2, 55, 75, 93, 16, 73, and 8. The 25 test views are sampled evenly from the testing images for evaluation. All images are downsampled $2\times$ to $400\times 400$ during the experiment.

### v3

@@ Abstract
Radiance fields have demonstrated impressive performance in synthesizing novel views from sparse input views, yet prevailing methods suffer from high training costs and slow inference speed. This paper introduces DNGaussian, a depth-regularized framework based on 3D Gaussian radiance fields, offering real-time and high-quality few-shot novel view synthesis at low costs. Our motivation stems from the highly efficient representation and surprising quality of the recent 3D Gaussian Splatting, despite it will encounter a geometry degradation when input views decrease. In the Gaussian radiance fields, we find this degradation in scene geometry primarily lined to the positioning of Gaussian primitives and can be mitigated by depth constraint. Consequently, we propose a Hard and Soft Depth Regularization to restore accurate scene geometry under coarse monocular depth supervision while maintaining a fine-grained color appearance. To further refine detailed geometry reshaping, we introduce Global-Local Depth Normalization, enhancing the focus on small local depth changes. Extensive experiments on LLFF, DTU, and Blender datasets demonstrate that DNGaussian outperforms state-of-the-art methods, achieving comparable or better results with significantly reduced memory cost, a 25×25$\times$25 × reduction in training time, and over 3000×3000$\times$3000 × faster rendering speed. Code is available at: https://github.com/Fictionarry/DNGaussian .
@@
$*{}^{*}$orresponding author: Xiao Bai (baixiao@buaa.edu.cn).
@@ 1 Introduction
Integrating the two proposed techniques, DNGaussian synthesizes views with competitive quality and superior details compared to state-of-the-art methods in multiple sparse-view settings on LLFF, Blender, and DTU datasets. This advantage is further enriched by substantially lower memory costs, 25×25$\times$25 × reduction of training time, and over 3000×3000$\times$3000 × faster rendering speed. The experiments also demonstrate our method’s universal ability to fit complex scenes, wide-ranging views, and multiple materials.
@@ 2 Related Work
Radiance Fields for Novel View Synthesis. Novel view synthesis aims to generate unseen views of the same object or scene from a set of given images [60, 1]. Neural Radiance Fields (NeRF) [25] uses a large MLP to represent 3D scenes and renders via volume rendering. However, its speed is slow both in training and inference. The following improvements mainly pursue either higher quality [2, 3] or efficiency [6, 26, 37, 12, 54, 21, 15], but hard to achieve both. The most recent unstructured radiance fields [7, 52, 18] utilize a set of primitives to represent scenes. Among them, 3D Gaussian Splatting [18] represents radiance fields by a set of anisotropic 3D Gaussians and renders with a differentiable splatting. This approach achieves great success in fast and high-quality reconstruction for complex real scenes. While this method excels with dense input views and has achieved success in various 3D tasks [23, 47, 39], its reconstruction with sparse view inputs remains an open problem. Also, issues such as how to apply additional constraints for improvement are still unsolved and worthy of discussion.
@@
Depth Supervision in Sparse-view Neural Fields. As a classic cue in many 3D vision tasks [46, 44, 41, 58, 43, 61], depth information has been widely used to supervise sparse-view neural fields. The first group [11, 31] is to extract accurate but sparse depth values from reliable point clouds, and the second [56, 14, 36, 40, 42] distills depth knowledge from current powerful monocular depth estimators [30, 29]. Considering point clouds are sparse and not available in many sparse-view cases, monocular depth shows its advantage in density and robustness for our tasks. To tackle the scale ambiguity and error of monocular depths, some previous works and concurrent sparse-view 3DGS methods have introduced various scale-invariant losses [56, 10, 36, 50, 63] including depth ranking loss [51, 42], however, all of which are not optimal for us. Firstly, flexible Gaussians are more sensitive to wrong depth cues, requiring extra designs for regularization. Also, these losses align the depth to a certain fixed global scale, which may ignore minor local depth changes. This overlook can lead to a noisy primitive distribution, particularly in regions with intricate textures. Besides, we notice an HDN loss [57] that can preserve details in monocular depth estimation. Nevertheless, it is also unsuitable as its reliance on multi-scale patches would bring long-distance errors and compromise geometric accuracy.
@@ 3.1 Preliminary for 3D Gaussian Splatting
Representation. 3D Gaussian splatting [18] represents 3D information with a set of 3D Gaussians. It computes pixel-wise color $\mathcal{C}$ with a set of 3D Gaussian primitives $\theta$, view pose $P$, and the camera parameter involving the center $o$.
@@
$\mathcal{G}_{i}(x)=e^{-\frac{1}{2}(x-\mu_{i})^{T}\Sigma_{i}^{-1}(x-\mu_{i})},$ (1)
@@
Rendering. 3D Gaussian Splatting utilizes a point-based rendering to compute the color $\mathcal{C}$ of pixel $x_{p}$ by blending $N$ ordered Gaussians overlapping the pixel:
@@
$\widetilde{\alpha}_{i}=\alpha_{i}\mathcal{G}^{proj}_{i}(x_{p}).$ (3)
Then, similar to NeRF, we can represent the pixel-wise depth $\mathcal{D}$ with the distance to the camera center $o$:
$\mathcal{D}(x_{p})=\sum_{i\in N}{||\mu_{i}-o||_{2}}\times\widetilde{\alpha}_{i}\prod_{j=1}^{i-1}(1-\widetilde{\alpha}_{j}).$ (4)
@@ 3.2 Depth Regularization for Gaussians
Hard Depth Regularization. To achieve the spatial reshaping of the Gaussian fields, we first propose a Hard Depth Regularization that encourages the movement of the nearest Gaussians, which are expected to compose surfaces but often cause noises and artifacts. Considering the predicted depth is rendered with the mixture of multiple Gaussians and reweighted by the cumulative product $\widetilde{\alpha}$, we manually apply a large opacity value $\tau$ to all Gaussians. Then, we render a “hard depth” that mainly consists of the nearest Gaussians on the ray shot from camera center $o$ and across the pixel $x_{p}$:
$\mathcal{D}_{hard}(x_{p})=\sum_{i\in N}{\tau(1-\tau)^{i-1}\mathcal{G}^{proj}_{i}(x_{p})||\mu_{i}-o||_{2}}.$ (5)
@@
Soft Depth Regularization. Only regularizing on “hard depth” is insufficient due to the absence of opacity optimization. We also expect to ensure the accuracy of the real rendered “soft depth”, otherwise, the surface may become semitransparent and cause hollowness. From this perspective, we additionally freeze the Gaussian center $\mu$ (denoted by $\check{\mu}$) to avoid negative influence caused by the center moving, and propose Soft Depth Regularization for the tuning of the opacity $\alpha$:
$\mathcal{D}_{soft}(x_{p})=\sum_{i\in N}{||\check{\mu_{i}}-o||_{2}}\times\widetilde{\alpha}_{i}\prod_{j=1}^{i-1}(1-\widetilde{\alpha}_{j}),$ (7)
@@ 3.3 Global-Local Depth Normalization
Local Depth Normalization. To solve this problem, we make the loss function refocus on small errors by introducing a patch-wise local normalization. Specifically, we cut a whole depth map into small patches and normalize the patch $\mathcal{P}$ of predicted depth and monocular depth with the mean value of 00 and standard deviation of near to $1$:
$\mathcal{D}^{LN}(x)=\frac{\mathcal{D}(x)-\text{mean}(\mathcal{D}(\mathcal{P}))}{\text{std}(\mathcal{D}(\mathcal{P}))+\epsilon},\quad\mathrm{s.t.}\ \ x\in\mathcal{P},$ (8)
@@
$\mathcal{D}^{GN}(x)=\frac{\mathcal{D}(x)-\text{mean}(\mathcal{D}(\mathcal{P}))}{\text{std}(\mathcal{D}_{I})},$ (9)
$\mathrm{s.t.}\ \ x\in\mathcal{P},$ $\mathcal{P}\subseteq\mathcal{I}.$
@@ 3.4 Training Details
$\mathcal{L}_{color}=\mathcal{L}_{1}(\hat{\mathcal{I}},\mathcal{I})+\lambda\mathcal{L}_{\mathrm{D-SSIM}}(\hat{\mathcal{I}},\mathcal{I}).$ (10)
@@
where $T$ stands for $hard$ or $soft$. In practice, we reserve an error tolerance for the L2 loss to relax the constraint. The full loss function is formulated by:
@@
MVSNeRF ft [5] Trained on DTU Fine-tuned per Scene 17.88 0.327 0.584 0.157 18.54 0.197 0.769 0.113
@@
Neural Color Renderer. 3D Gaussian Splatting stores color via spherical harmonics, however, it is easy to overfit with only sparse views. To relieve this problem, we take a grid encoder and an MLP as the Neural Color Renderer to predict color for each primitive (Figure 3). During inference, we store the intermediate result and only calculate the last MLP layers to merge view direction for acceleration.
@@ 4.1 Setups
Datasets. we conduct our experiment on three datasets: the NeRF Blender Synthetic dataset (Blender) [25], the DTU dataset [17], and the LLFF dataset [24]. We follow the setting used in previous works [27, 53, 42] with the same split of DTU and LLFF to train the model on 3 views and test on another set of images. To erase the noises in the background and focus on the target object, we apply the same object masks as previous works [27] for DTU at evaluation. For Blender, we follow DietNeRF [16] and FreeNeRF [53] to train with the same 8 views and test on 25 unseen images. Aligned with the baselines, downsampling rates of $8$, $4$, and $2$ are applied to LLFF, DTU, and Blender. Following previous sparse-view settings, the camera poses are assumed to be known via calibration or other ways.
@@ 4.2 Comparison
Efficiency. We further conduct an efficiency study on the LLFF 3-view setting with RTX 3090 Ti GPUs to explore the performance of current SOTA baselines [42, 53] with limited GPU memories of 24GB/12GB, and training time of 1.0h/0.5h, as shown in Table 3. The top row of each group represents the default setting of the corresponding baseline, where the training time is measured by us with the same number of iterations on a single GPU. While both FreeNeRF and SparseNeRF perform worse under strict resource limitations, our method shows huge advantages in efficiency, which achieves remarkable accelerations of 25×25$\times$25 × on training time and over 3000×3000$\times$3000 × on FPS, while synthesizing competitive quality novel views. Given the necessity for per-scene optimization and rapid visualization, our high efficiency holds significant value for practical applications.
@@ 5 Conclusion
Acknowledgements. In this work, we are supported by the National Natural Science Foundation of China 62276016, 62372029. Lin Gu is supported by JST Moonshot R&D Grant Number JPMJMS2011 Japan.
@@ A.2 Ablation Study on Neural Color Renderer
In this work, we replace the spherical harmonic (SH) of 3D Gaussian Splatting with a neural color renderer to represent the direction-variant color. To better illustrate the function of this module, we compare it to the original SH function with different degrees in the LLFF dataset with 3 training views. The results are in Table 7 and Figure 11. The SH function is easy to overfit in the sparse-view situation and results in some strange colors during view changing. This may be caused by the independence of each primitive which leads to a lack of regional consistency. After introducing the neural color renderer, the problem has been relieved. By storing the intermediate result and only calculating the latest two MLP layers, we can maintain a fast rendering speed competitive to SH as well.
@@ B.1 Implementations
Metrics. Following previous methods [27, 53], we utilize the “structural_similarity” API from scikit-image  [^3] https://scikit-image.org/docs/stable/api/skimage.metrics.html to compute the SSIM score, and use the implementation with a pre-trained VGG model to calculate the LPIPS score.
@@ B.2 Datasets
LLFF. The LLFF dataset [24] contains 8 forward-facing scenes in total. Following [27, 53, 42], we take every 8-th image as the novel views for testing. The input views are evenly sampled across the remaining views. Images are downsampled 8×8$\times$8 × to the resolution of $378\times 504$. In practice, we ignore the distortion of the original images.
DTU. The DTU dataset [17] consists of 124 object-centric scenes captured by a set of fixed cameras. We follow [27, 53, 42] to evaluate models directly on the 15 scenes with the scan IDs of 8, 21, 30, 31, 34, 38, 40, 41, 45, 55, 63, 82, 103, 110, and 114. In each scan, the images with the following IDs of 25, 22, and 28 are used as the input views in our 3-view setting. The test set consists of images with IDs of 1, 2, 9, 10, 11, 12, 14, 15, 23, 24, 26, 27, 29, 30, 31, 32, 33, 34, 35, 41, 42, 43, 45, 46 and 47 for evaluation. The images are downsampled 4×4$\times$4 ×. In particular, we use the undistorted images from COLMAP to eliminate the negative impact of unerased lens distortion.
Blender. We follow the data split used in [16, 53] for the Blender dataset [11]. The 8 input views are selected from the training images, with IDs 26, 86, 2, 55, 75, 93, 16, 73, and 8. The 25 test views are sampled evenly from the testing images for evaluation. All images are downsampled 2×2$\times$2 × to $400\times 400$ during the experiment.